# Critical point representation of the mutual information in the sparse stochastic block model

### v1

@@ 4 Cavity computation
In this section we perform the cavity computation that gives rise to the functional (1.26) and leads to the critical point representation in Theorem 1.1. The idea behind the cavity computation is to determine the effect of perturbing the Hamiltonian (3.9) for the system of size $N$ by a single additional spin. To emphasize this additional spin, we write $\rho:=(\epsilon,\sigma)\in\Sigma_{N+1}$ for $\sigma\in\Sigma_{N}$ and $\epsilon\in\{-1,+1\}$. Similarly, we write $\rho^{*}:=(\epsilon^{*},\sigma^{*})\in\Sigma_{N+1}$ for $\sigma^{*}\in\Sigma_{N}$ and $\epsilon^{*}\in\Sigma_{1}$ to denote the ground truth sampled from $P_{N+1}^{*}$. Given a pair $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ and a perturbation parameter $\lambda=(\lambda_{k})_{k\geq 0}$, we will compare the systems with Hamiltonians $H_{N+1}^{t,\mu}(\epsilon,\sigma,\lambda)$ and $H_{N}^{t,\mu}(\sigma,\lambda)$. This comparison will be done at the level of the free energies, and the re-scaled difference between $\overline{F}_{N+1}(t,\mu,\lambda)$ and $\overline{F}_{N}(t,\mu,\lambda)$ will be expressed as the difference between the modified Gibbs averages of two quantities that we may call the “cavity fields”. We now describe the modified Hamiltonian associated with this modified Gibbs average as well as the two cavity fields.
@@
$\mathbb{P}\big{\{}{G^{\prime}}^{k}_{i}=1\mid\epsilon^{*},\sigma^{*}\big{\}}:=\frac{c+\Delta\epsilon^{*}\sigma_{i}^{*}}{N+1}.$ (4.10)
Given a collection of random indices $(\ell_{k})_{k\geq 1}$ sampled uniformly at random from $\{1,\ldots,N\}$, independently of the other random variables, we define the time-dependent cavity field $z_{N}^{t}(\epsilon,\sigma)$ on $\Sigma_{N+1}$ by
$z_{N}^{t}(\epsilon,\sigma):=\sum_{k\leq\Pi_{t}^{\prime}}\log\bigg{[}\big{(}c+\Delta\epsilon\sigma_{\ell_{k}}\big{)}^{{G^{\prime}}^{k}_{\ell_{k}}}\Big{(}1-\frac{c+\Delta\epsilon\sigma_{\ell_{k}}}{N+1}\Big{)}^{1-{G^{\prime}}^{k}_{\ell_{k}}}\bigg{]}.$ (4.11)
Similarly, for each $\mu\in\mathcal{M}_{+}$, we define the measure-dependent cavity field $z_{N}^{\mu}(\epsilon)$ on $\Sigma_{1}$ by
$z_{N}^{\mu}(\epsilon):=\sum_{x\in\Lambda_{i}(N\mu)}\log\bigg{[}\big{(}c+\Delta\epsilon x\big{)}^{{G^{\prime}}^{x}}\Big{(}1-\frac{c+\Delta\epsilon x}{N+1}\Big{)}^{1-{G^{\prime}}^{x}}\bigg{]},$ (4.12)
@@
$\mathbb{P}\big{\{}G^{\prime x}=1\mid\epsilon^{*},x\big{\}}:=\frac{c+\Delta\epsilon^{*}x}{N+1}.$ (4.13)
For each pair $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$, the cavity field $z_{N}^{t,\mu}(\epsilon,\sigma)$ is the sum of the time-dependent and measure-dependent cavity fields,
$z_{N}^{t,\mu}(\epsilon,\sigma):=z_{N}^{t}(\epsilon,\sigma)+z_{N}^{\mu}(\epsilon).$ (4.14)
@@
$A_{N}(t,\mu,\lambda):=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}^{t,\mu}(\epsilon,\sigma)\rangle^{\prime}\,\mathrm{d}P^{*}(\epsilon)-\operatorname{\mathbb{E}}\log\big{\langle}\exp y_{N}^{t}(\sigma)\big{\rangle}^{\prime}.$ (4.17)
@@
For simplicity, we will ignore the perturbation Hamiltonians (3.3) and (3.7); however, the choices (3.5)-(3.6) of the sequences $(\epsilon_{N})_{N\geq 1}$ and $(s_{N})_{N\geq 1}$ can be used to show that the error incurred by replacing the perturbation Hamiltonian $H_{N+1}^{\mathrm{gauss}}(\sigma,\lambda_{0})+H_{N+1}^{\mathrm{exp}}(\sigma,\lambda)$ for the system of size $N+1$ by that for the system of size $N$ can be absorbed into the error term on the right side of (4.19). The proof proceeds in two steps. First we relate the time-dependent Hamiltonian (1.15) for the system of size $N+1$ to that for the system of size $N$, and then we do the same for the measure-dependent Hamiltonian (1.18).
@@
$H_{N+1}^{t}(\rho)\stackrel{{\scriptstyle d}}{{\approx}}\mathcal{H}_{N}^{t}(\sigma)+z_{N}^{t}(\epsilon,\sigma)$ (4.20)
@@
$H_{N+1}^{\mu}(\rho)\stackrel{{\scriptstyle d}}{{\approx}}\mathcal{H}_{N}^{\mu}(\sigma)+z_{N}^{\mu}(\epsilon,\sigma)$ (4.22)
@@
$A_{N}^{1}(t,\mu,\lambda^{N}):=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}^{t,\mu}(\epsilon,\sigma)\rangle^{\prime}\,\mathrm{d}P^{*}(\epsilon)\quad\text{and}\quad A_{N}^{2}(t,\mu,\lambda^{N}):=\operatorname{\mathbb{E}}\log\langle\exp y_{N}^{t}(\sigma)\rangle^{\prime}$
@@
$A_{N}^{1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}(\epsilon,\sigma^{1})\rangle^{\prime}\,\mathrm{d}P^{*}(\epsilon)$
@@
$z_{N}(\epsilon,\sigma^{1}):=\sum_{k\leq\Pi_{t}^{\prime}}\log\bigg{[}\big{(}c+\Delta\epsilon\sigma_{\ell_{k}}^{1}\big{)}^{{G^{\prime}}^{k}_{\ell_{k}}}\Big{(}1-\frac{c+\Delta\epsilon\sigma_{\ell_{k}}^{1}}{N+1}\Big{)}^{1-{G^{\prime}}^{k}_{\ell_{k}}}\bigg{]}+z_{N}^{\mu}(\epsilon),$
@@
$\mathbb{P}\big{\{}{G^{\prime}}^{k}_{i}=1\mid\epsilon^{*},\sigma^{2}\big{\}}:=\frac{c+\Delta\epsilon^{*}\sigma_{i}^{2}}{N+1}.$
The Poisson coloring theorem implies that the number of terms in the first sum defining $z_{N}(\epsilon,\sigma^{1})$ that have $\smash{{G^{\prime}}^{k}_{\ell_{k}}}$ equal to zero is Poisson with mean
$tN\cdot\bigg{(}1-\frac{c+\Delta\epsilon^{*}\sigma_{\ell_{k}}^{2}}{N+1}\bigg{)}=tN+\operatorname{\operatorname{\mathcal{O}}}(1).$
Together with a Taylor expansion of the logarithm and the law of large numbers, this implies that the contribution of the terms in the first term defining $z_{N}(\epsilon,\sigma^{1})$ that have $\smash{{G^{\prime}}^{k}_{\ell_{k}}}$ equal to zero is
$-\sum_{k\leq tN}\frac{c+\Delta\epsilon\sigma^{1}_{\ell_{k}}}{N}+\operatorname{\operatorname{\mathcal{O}}}\big{(}N^{-1}\big{)}=-ct-\Delta t\epsilon\overline{m}+\operatorname{\operatorname{\mathcal{O}}}\big{(}N^{-1}\big{)}.$
A similar argument shows that the contribution of the terms in $z_{N}^{\mu}(\epsilon)$ that have $\smash{{G^{\prime}}^{x}}$ equal to zero is $-cs-\Delta s\epsilon\operatorname{\mathbb{E}}x_{1}+\operatorname{\operatorname{\mathcal{O}}}(N^{-1})$. If we define
$z_{N}^{\prime}(\epsilon,\sigma^{1}):=\sum_{k\leq\Pi_{t}^{\prime}}\log\big{(}c+\Delta\epsilon\sigma_{\ell_{k}}^{1}\big{)}^{{G^{\prime}}^{k}_{\ell_{k}}}+\sum_{k\leq\Pi^{\prime}_{s}}\log\big{(}c+\Delta\epsilon x_{k}\big{)}^{{G^{\prime}}_{k}^{x}},$
@@
$\mathbb{P}\big{\{}{G^{\prime}}_{k}^{x}=1\mid\epsilon^{*},x_{k}\big{\}}:=\frac{c+\Delta\epsilon^{*}x_{k}}{N+1},$
@@
$A_{N}^{1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\langle\exp z_{N}^{\prime}(\epsilon,\sigma^{1})\rangle^{\prime}\,\mathrm{d}P^{*}(\epsilon)-c(t+s)+\operatorname{\operatorname{\mathcal{O}}}\big{(}N^{-1}\big{)}.$ (4.26)
@@
$\binom{\Pi_{t}^{\prime}}{r}\binom{\Pi_{s}^{\prime}}{q}\prod_{i=1}^{r}$ $\frac{c+\Delta\epsilon^{*}\sigma_{i}^{2}}{N+1}\prod_{i=r+1}^{\Pi_{t}^{\prime}}\bigg{(}1-\frac{c+\Delta\epsilon^{*}\sigma_{i}^{2}}{N+1}\bigg{)}\prod_{j=1}^{q}\frac{c+\Delta\epsilon^{*}x_{j}}{N+1}\prod_{j=q+1}^{\Pi_{s}^{\prime}}\bigg{(}1-\frac{c+\Delta\epsilon^{*}x_{j}}{N+1}\bigg{)}$
$\qquad\log\bigg{\langle}\int_{\Sigma_{1}}\exp(-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{i=1}^{r}(c+\Delta\epsilon\sigma^{1}_{i})\prod_{j=1}^{q}(c+\Delta\epsilon x_{j})\,\mathrm{d}P^{*}(\epsilon)\bigg{\rangle}^{\prime}$
with respect to the randomness of $\sigma^{2}$, $(x_{k})_{k\geq 1}$, $\Pi_{t}^{\prime}$, $\Pi_{s}^{\prime}$, and $\epsilon^{*}$. Since the expression in the second line of this display grows at most linearly in $r$ and $q$, the concentration of $\Pi_{t}^{\prime}$ and $\Pi_{s}^{\prime}$ about $tN$ and $sN$, respectively, and the rapid decay of the binomial coefficients allow us to truncate both sums at a large enough integer $M$ by incurring an error of at most $\delta$ uniformly over $N$. Together with the approximations
$\prod_{i=r+1}^{\Pi_{t}^{\prime}}\bigg{(}1-\frac{c+\Delta\epsilon^{*}\sigma_{i}^{1}}{N+1}\bigg{)}$ $=\exp(-tc-\Delta t\epsilon^{*}\overline{m})+o_{N}(1),$
$\prod_{j=q+1}^{\Pi_{s}^{\prime}}\bigg{(}1-\frac{c+\Delta\epsilon^{*}x_{j}}{N+1}\bigg{)}$ $=\exp(-sc-\Delta s\epsilon^{*}\operatorname{\mathbb{E}}x_{1})+o_{N}(1),$
@@
$\operatorname{\mathbb{E}}\sum_{r\leq\Pi_{t}^{\prime}\wedge M}\sum_{q\leq\Pi_{s}^{\prime}\wedge M}\binom{\Pi_{t}^{\prime}}{r}$ $\binom{\Pi_{s}^{\prime}}{q}\frac{e^{-c(t+s)-\Delta\epsilon^{*}(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}}{(N+1)^{r+q}}\bigg{\langle}\prod_{i=1}^{r}(c+\Delta\epsilon^{*}\sigma_{i}^{2})\prod_{j=1}^{q}(c+\Delta\epsilon^{*}x_{j})$
$\log\bigg{\langle}\int_{\Sigma_{1}}e^{-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}\prod_{i=1}^{r}(c+\Delta\epsilon\sigma^{1}_{i})\prod_{j=1}^{q}(c+\Delta\epsilon x_{j})\,\mathrm{d}P^{*}(\epsilon)\bigg{\rangle}^{\prime}\bigg{\rangle}^{\prime}.$
@@
$A_{N}^{1,1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\exp z_{N}^{\prime}(\epsilon)\,\mathrm{d}P^{*}(\epsilon)+o_{N}(1),$
@@
$z_{N}^{\prime}(\epsilon):=\sum_{k\leq\Pi_{t}^{\prime}}\log(c+\Delta\epsilon y_{k})^{{G^{\prime}}^{y}_{k}}+\sum_{k\leq\Pi_{s}^{\prime}}\log(c+\Delta\epsilon x_{k})^{{G^{\prime}}^{x}_{k}},$ (4.27)
@@
$A_{N}^{1,1}$ $=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{k\leq\Pi_{t}^{\prime}}(c+\Delta\epsilon y_{k})^{{G^{\prime}}_{k}^{y}}\prod_{k\leq\Pi_{s}^{\prime}}(c+\Delta\epsilon x_{k})^{{G^{\prime}}_{k}^{x}}\,\mathrm{d}P^{*}(\epsilon)+o_{N}(1)$
$=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{\ell\leq K}\prod_{k\in I_{t}^{\ell}\cup I_{s}^{\ell}}(c+\Delta\epsilon a_{\ell})\,\mathrm{d}P^{*}(\epsilon)+o_{N}(1).$
@@
$\operatorname{\mathbb{E}}\Pi_{t}^{\prime}\mathbb{P}\{{G^{\prime}}_{1}^{y}=1,y_{1}=a_{\ell}\}+\operatorname{\mathbb{E}}\Pi_{s}^{\prime}\mathbb{P}\{{G^{\prime}}_{1}^{x}=1,x_{1}=a_{\ell}\}$ $=sp_{\ell}(c+\Delta\epsilon^{*}a_{\ell})+tq_{\ell}(c+\Delta\epsilon^{*}a_{\ell})+o_{N}(1)$
$=(c+\Delta\epsilon^{*}a_{\ell})(\mu+t\nu)(a_{\ell})+o_{N}(1).$
@@
$A_{N}^{1,1}$ $=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\epsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{x\in\Pi_{*}(\mu+t\nu)}(c+\Delta\epsilon x)\,\mathrm{d}P^{*}(\epsilon)+o_{N}(1)$
@@
where $\Pi_{*}(\mu+t\nu)$ is a Poisson point process with mean measure $(c+\Delta\epsilon^{*}x)\,\mathrm{d}(\mu+t\nu)(x)$. Extending this result to the case when $\mu$ and $\nu$ are arbitrary can be done through a continuity argument leveraging Proposition 2.1. Combining this with (4.26) establishes the first equality in (4.25).
@@
$-\sum_{k\leq tN^{2}/2}\frac{c+\Delta\sigma^{1}_{\ell_{k}}\sigma^{1}_{m_{k}}}{N(N+1)}+\operatorname{\operatorname{\mathcal{O}}}\big{(}N^{-1}\big{)}=-\frac{t}{2}\big{(}c+\Delta\overline{m}^{2}\big{)}+\operatorname{\operatorname{\mathcal{O}}}\big{(}N^{-1}\big{)}.$
@@
[…]
@@
[…]

### v2

@@ 4 Cavity computation
In this section we perform the cavity computation that gives rise to the functional (1.26) and leads to the critical point representation in Theorem 1.1. The idea behind the cavity computation is to determine the effect of perturbing the Hamiltonian (3.9) for the system of size $N$ by a single additional spin. To emphasize this additional spin, we write $\rho:=(\varepsilon,\sigma)\in\Sigma_{N+1}$ for $\sigma\in\Sigma_{N}$ and $\varepsilon\in\{-1,+1\}$. Similarly, we write $\rho^{*}:=(\varepsilon^{*},\sigma^{*})\in\Sigma_{N+1}$ for $\sigma^{*}\in\Sigma_{N}$ and $\varepsilon^{*}\in\Sigma_{1}$ to denote the ground truth sampled from $P_{N+1}^{*}$. Given a pair $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ and a perturbation parameter $\lambda=(\lambda_{k})_{k\geq 0}$, we will compare the systems with Hamiltonians $H_{N+1}^{t,\mu}(\varepsilon,\sigma,\lambda)$ and $H_{N}^{t,\mu}(\sigma,\lambda)$. This comparison will be done at the level of the free energies, and the re-scaled difference between $\overline{F}_{N+1}(t,\mu,\lambda)$ and $\overline{F}_{N}(t,\mu,\lambda)$ will be expressed as the difference between the modified Gibbs averages of two quantities that we may call the “cavity fields”. We now describe the modified Hamiltonian associated with this modified Gibbs average as well as the two cavity fields.
@@
$\mathbb{P}\big\{{G^{\prime}}^{k}_{i}=1\mid\varepsilon^{*},\sigma^{*}\big\}:=\frac{c+\Delta\varepsilon^{*}\sigma_{i}^{*}}{N+1}.$ (4.10)
Given a collection of random indices $(\ell_{k})_{k\geq 1}$ sampled uniformly at random from $\{1,\ldots,N\}$, independently of the other random variables, we define the time-dependent cavity field $z_{N}^{t}(\varepsilon,\sigma)$ on $\Sigma_{N+1}$ by
$z_{N}^{t}(\varepsilon,\sigma):=\sum_{k\leq\Pi_{t}^{\prime}}\log\bigg[\big(c+\Delta\varepsilon\sigma_{\ell_{k}}\big)^{{G^{\prime}}^{k}_{\ell_{k}}}\Big(1-\frac{c+\Delta\varepsilon\sigma_{\ell_{k}}}{N+1}\Big)^{1-{G^{\prime}}^{k}_{\ell_{k}}}\bigg].$ (4.11)
Similarly, for each $\mu\in\mathcal{M}_{+}$, we define the measure-dependent cavity field $z_{N}^{\mu}(\varepsilon)$ on $\Sigma_{1}$ by
$z_{N}^{\mu}(\varepsilon):=\sum_{x\in\Lambda_{i}(N\mu)}\log\bigg[\big(c+\Delta\varepsilon x\big)^{{G^{\prime}}^{x}}\Big(1-\frac{c+\Delta\varepsilon x}{N+1}\Big)^{1-{G^{\prime}}^{x}}\bigg],$ (4.12)
@@
$\mathbb{P}\big\{G^{\prime x}=1\mid\varepsilon^{*},x\big\}:=\frac{c+\Delta\varepsilon^{*}x}{N+1}.$ (4.13)
For each pair $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$, the cavity field $z_{N}^{t,\mu}(\varepsilon,\sigma)$ is the sum of the time-dependent and measure-dependent cavity fields,
$z_{N}^{t,\mu}(\varepsilon,\sigma):=z_{N}^{t}(\varepsilon,\sigma)+z_{N}^{\mu}(\varepsilon).$ (4.14)
@@
$A_{N}(t,\mu,\lambda):=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}^{t,\mu}(\varepsilon,\sigma)\rangle^{\prime}\,\mathrm{d}P^{*}(\varepsilon)-\operatorname{\mathbb{E}}\log\big\langle\exp y_{N}^{t}(\sigma)\big\rangle^{\prime}.$ (4.17)
@@
For simplicity, we will ignore the perturbation Hamiltonians (3.3) and (3.7); however, the choices (3.5)-(3.6) of the sequences $(\varepsilon_{N})_{N\geq 1}$ and $(s_{N})_{N\geq 1}$ can be used to show that the error incurred by replacing the perturbation Hamiltonian $H_{N+1}^{\mathrm{gauss}}(\sigma,\lambda_{0})+H_{N+1}^{\mathrm{exp}}(\sigma,\lambda)$ for the system of size $N+1$ by that for the system of size $N$ can be absorbed into the error term on the right side of (4.19). The proof proceeds in two steps. First we relate the time-dependent Hamiltonian (1.15) for the system of size $N+1$ to that for the system of size $N$, and then we do the same for the measure-dependent Hamiltonian (1.18).
@@
$H_{N+1}^{t}(\rho)\stackrel{{\scriptstyle d}}{{\approx}}\mathcal{H}_{N}^{t}(\sigma)+z_{N}^{t}(\varepsilon,\sigma)$ (4.20)
@@
$H_{N+1}^{\mu}(\rho)\stackrel{{\scriptstyle d}}{{\approx}}\mathcal{H}_{N}^{\mu}(\sigma)+z_{N}^{\mu}(\varepsilon,\sigma)$ (4.22)
@@
$A_{N}^{1}(t,\mu,\lambda^{N}):=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}^{t,\mu}(\varepsilon,\sigma)\rangle^{\prime}\,\mathrm{d}P^{*}(\varepsilon)\quad\text{and}\quad A_{N}^{2}(t,\mu,\lambda^{N}):=\operatorname{\mathbb{E}}\log\langle\exp y_{N}^{t}(\sigma)\rangle^{\prime}$
@@
$A_{N}^{1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\langle\exp z_{N}(\varepsilon,\sigma^{1})\rangle^{\prime}\,\mathrm{d}P^{*}(\varepsilon)$
@@
$z_{N}(\varepsilon,\sigma^{1}):=\sum_{k\leq\Pi_{t}^{\prime}}\log\bigg[\big(c+\Delta\varepsilon\sigma_{\ell_{k}}^{1}\big)^{{G^{\prime}}^{k}_{\ell_{k}}}\Big(1-\frac{c+\Delta\varepsilon\sigma_{\ell_{k}}^{1}}{N+1}\Big)^{1-{G^{\prime}}^{k}_{\ell_{k}}}\bigg]+z_{N}^{\mu}(\varepsilon),$
@@
$\mathbb{P}\big\{{G^{\prime}}^{k}_{i}=1\mid\varepsilon^{*},\sigma^{2}\big\}:=\frac{c+\Delta\varepsilon^{*}\sigma_{i}^{2}}{N+1}.$
The Poisson coloring theorem implies that the number of terms in the first sum defining $z_{N}(\varepsilon,\sigma^{1})$ that have $\smash{{G^{\prime}}^{k}_{\ell_{k}}}$ equal to zero is Poisson with mean
$tN\cdot\bigg(1-\frac{c+\Delta\varepsilon^{*}\sigma_{\ell_{k}}^{2}}{N+1}\bigg)=tN+\operatorname{\operatorname{\mathcal{O}}}(1).$
Together with a Taylor expansion of the logarithm and the law of large numbers, this implies that the contribution of the terms in the first term defining $z_{N}(\varepsilon,\sigma^{1})$ that have $\smash{{G^{\prime}}^{k}_{\ell_{k}}}$ equal to zero is
$\sum_{k\leq tN}\frac{c+\Delta\varepsilon\sigma^{1}_{\ell_{k}}}{N}+\operatorname{\operatorname{\mathcal{O}}}\big(N^{-1}\big)=-ct-\Delta t\varepsilon\overline{m}+\operatorname{\operatorname{\mathcal{O}}}\big(N^{-1}\big).$
A similar argument shows that the contribution of the terms in $z_{N}^{\mu}(\varepsilon)$ that have $\smash{{G^{\prime}}^{x}}$ equal to zero is $cs-\Delta s\varepsilon\operatorname{\mathbb{E}}x_{1}+\operatorname{\operatorname{\mathcal{O}}}(N^{-1})$. If we define
$z_{N}^{\prime}(\varepsilon,\sigma^{1}):=\sum_{k\leq\Pi_{t}^{\prime}}\log\big(c+\Delta\varepsilon\sigma_{\ell_{k}}^{1}\big)^{{G^{\prime}}^{k}_{\ell_{k}}}+\sum_{k\leq\Pi^{\prime}_{s}}\log\big(c+\Delta\varepsilon x_{k}\big)^{{G^{\prime}}_{k}^{x}},$
@@
$\mathbb{P}\big\{{G^{\prime}}_{k}^{x}=1\mid\varepsilon^{*},x_{k}\big\}:=\frac{c+\Delta\varepsilon^{*}x_{k}}{N+1},$
@@
$A_{N}^{1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\langle\exp z_{N}^{\prime}(\varepsilon,\sigma^{1})\rangle^{\prime}\,\mathrm{d}P^{*}(\varepsilon)-c(t+s)+\operatorname{\operatorname{\mathcal{O}}}\big(N^{-1}\big).$ (4.26)
@@
$\binom{\Pi_{t}^{\prime}}{r}\binom{\Pi_{s}^{\prime}}{q}\prod_{i=1}^{r}$ $\frac{c+\Delta\varepsilon^{*}\sigma_{i}^{2}}{N+1}\prod_{i=r+1}^{\Pi_{t}^{\prime}}\bigg(1-\frac{c+\Delta\varepsilon^{*}\sigma_{i}^{2}}{N+1}\bigg)\prod_{j=1}^{q}\frac{c+\Delta\varepsilon^{*}x_{j}}{N+1}\prod_{j=q+1}^{\Pi_{s}^{\prime}}\bigg(1-\frac{c+\Delta\varepsilon^{*}x_{j}}{N+1}\bigg)$
$\qquad\log\bigg\langle\int_{\Sigma_{1}}\exp(-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{i=1}^{r}(c+\Delta\varepsilon\sigma^{1}_{i})\prod_{j=1}^{q}(c+\Delta\varepsilon x_{j})\,\mathrm{d}P^{*}(\varepsilon)\bigg\rangle^{\prime}$
with respect to the randomness of $\sigma^{2}$, $(x_{k})_{k\geq 1}$, $\Pi_{t}^{\prime}$, $\Pi_{s}^{\prime}$, and $\varepsilon^{*}$. Since the expression in the second line of this display grows at most linearly in $r$ and $q$, the concentration of $\Pi_{t}^{\prime}$ and $\Pi_{s}^{\prime}$ about $tN$ and $sN$, respectively, and the rapid decay of the binomial coefficients allow us to truncate both sums at a large enough integer $M$ by incurring an error of at most $\delta$ uniformly over $N$. Together with the approximations
$\prod_{i=r+1}^{\Pi_{t}^{\prime}}\bigg(1-\frac{c+\Delta\varepsilon^{*}\sigma_{i}^{1}}{N+1}\bigg)$ $=\exp(-tc-\Delta t\varepsilon^{*}\overline{m})+o_{N}(1),$
$\prod_{j=q+1}^{\Pi_{s}^{\prime}}\bigg(1-\frac{c+\Delta\varepsilon^{*}x_{j}}{N+1}\bigg)$ $=\exp(-sc-\Delta s\varepsilon^{*}\operatorname{\mathbb{E}}x_{1})+o_{N}(1),$
@@
$\operatorname{\mathbb{E}}\sum_{r\leq\Pi_{t}^{\prime}\wedge M}\sum_{q\leq\Pi_{s}^{\prime}\wedge M}\binom{\Pi_{t}^{\prime}}{r}$ $\binom{\Pi_{s}^{\prime}}{q}\frac{e^{-c(t+s)-\Delta\varepsilon^{*}(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}}{(N+1)^{r+q}}\bigg\langle\prod_{i=1}^{r}(c+\Delta\varepsilon^{*}\sigma_{i}^{2})\prod_{j=1}^{q}(c+\Delta\varepsilon^{*}x_{j})$
$\log\bigg\langle\int_{\Sigma_{1}}e^{-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1})}\prod_{i=1}^{r}(c+\Delta\varepsilon\sigma^{1}_{i})\prod_{j=1}^{q}(c+\Delta\varepsilon x_{j})\,\mathrm{d}P^{*}(\varepsilon)\bigg\rangle^{\prime}\bigg\rangle^{\prime}.$
@@
$A_{N}^{1,1}=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\exp z_{N}^{\prime}(\varepsilon)\,\mathrm{d}P^{*}(\varepsilon)+o_{N}(1),$
@@
$z_{N}^{\prime}(\varepsilon):=\sum_{k\leq\Pi_{t}^{\prime}}\log(c+\Delta\varepsilon y_{k})^{{G^{\prime}}^{y}_{k}}+\sum_{k\leq\Pi_{s}^{\prime}}\log(c+\Delta\varepsilon x_{k})^{{G^{\prime}}^{x}_{k}},$ (4.27)
@@
$A_{N}^{1,1}$ $=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{k\leq\Pi_{t}^{\prime}}(c+\Delta\varepsilon y_{k})^{{G^{\prime}}_{k}^{y}}\prod_{k\leq\Pi_{s}^{\prime}}(c+\Delta\varepsilon x_{k})^{{G^{\prime}}_{k}^{x}}\,\mathrm{d}P^{*}(\varepsilon)+o_{N}(1)$
$=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{\ell\leq K}\prod_{k\in I_{t}^{\ell}\cup I_{s}^{\ell}}(c+\Delta\varepsilon a_{\ell})\,\mathrm{d}P^{*}(\varepsilon)+o_{N}(1).$
@@
$\operatorname{\mathbb{E}}\Pi_{t}^{\prime}\mathbb{P}\{{G^{\prime}}_{1}^{y}=1,y_{1}=a_{\ell}\}+\operatorname{\mathbb{E}}\Pi_{s}^{\prime}\mathbb{P}\{{G^{\prime}}_{1}^{x}=1,x_{1}=a_{\ell}\}$ $=sp_{\ell}(c+\Delta\varepsilon^{*}a_{\ell})+tq_{\ell}(c+\Delta\varepsilon^{*}a_{\ell})+o_{N}(1)$
$=(c+\Delta\varepsilon^{*}a_{\ell})(\mu+t\nu)(a_{\ell})+o_{N}(1).$
@@
$A_{N}^{1,1}$ $=\operatorname{\mathbb{E}}\log\int_{\Sigma_{1}}\exp(-\Delta\varepsilon(t\overline{m}+s\operatorname{\mathbb{E}}x_{1}))\prod_{x\in\Pi_{*}(\mu+t\nu)}(c+\Delta\varepsilon x)\,\mathrm{d}P^{*}(\varepsilon)+o_{N}(1)$
@@
where $\Pi_{*}(\mu+t\nu)$ is a Poisson point process with mean measure $(c+\Delta\varepsilon^{*}x)\,\mathrm{d}(\mu+t\nu)(x)$. Extending this result to the case when $\mu$ and $\nu$ are arbitrary can be done through a continuity argument leveraging Proposition 2.1. Combining this with (4.26) establishes the first equality in (4.25).
@@
$\sum_{k\leq tN^{2}/2}\frac{c+\Delta\sigma^{1}_{\ell_{k}}\sigma^{1}_{m_{k}}}{N(N+1)}+\operatorname{\operatorname{\mathcal{O}}}\big(N^{-1}\big)=-\frac{t}{2}\big(c+\Delta\overline{m}^{2}\big)+\operatorname{\operatorname{\mathcal{O}}}\big(N^{-1}\big).$
@@
Combining Lemma 3.2 with the arguments in [43, Lemma 3.3], it is possible to find a sequence of perturbation parameters $(\lambda^{N})_{N\geq 1}$ such that the multioverlap concentration (3.15) for the perturbed Hamiltonian (4.6) holds for all $n\geq 1$. Remembering that the perturbation parameters do not affect the asymptotic behavior of the free energy by Lemma 3.1 and that the $\liminf$ of a Cesàro sum is bounded from below by the $\liminf$ of its general term, we see that
@@
Invoking Lemma 4.2 and passing to a further subsequence along which $(A_{N}(t,\mu,\lambda^{N}))_{N\geq 1}$ converges to its $\liminf$ if necessary, gives a sequence $\smash{(N_{k}^{-},\lambda^{N_{k}^{-}})_{k\geq 1}}$ and a measure $\nu^{-}\in\mathcal{M}_{p}$ such that the asymptotic spin array associated with the perturbation parameters $\smash{(\lambda^{N_{k}^{-}})_{k\geq 1}}$ is generated by $\nu^{-}$ according to (S1)-(S2) and